OpenStudy (anonymous):

Circle P has a radius of 8 units with center P at (-2, 1). Which equation defines circle P? (x + 2)2 + (y – 1)2 = 8 (x – 2)2 + (y + 1)2 = 8 (x + 2)2 + (y – 1)2 = 64 (x – 2)2 + (y + 1)2 = 64

OpenStudy (cwrw238):

equation of a circle is (x - a)^2 + (y - b)^2 = r^2 where (a,b) is center and r = radius so if the x coordinate of circle = -2 then first term is (x + 2)^2

OpenStudy (anonymous):

is it (x + 2)2 + (y – 1)2 = 8 ?

OpenStudy (cwrw238):

not quite - its r^2

OpenStudy (anonymous):

(x + 2)2 + (y – 1)2 = 64 ?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

thanks !